# Computational Complexity of Sufficiency in Decision Problems

### v2

@@ Abstract
We characterize the computational complexity of coordinate sufficiency in decision problems within the formal model. Given action set $A$, state space $S=X_{1}\times\cdots\times X_{n}$, and utility $u:A\times S\to\mathbb{R}$, a coordinate set $I$ is sufficient if $s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime})$.
The landscape in the formal model:
  • General case: SUFFICIENCY-CHECK is coNP-complete; ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete.
  • Tractable cases: Polynomial-time for bounded action sets under the explicit-state encoding; separable utilities ($u=f+g$) under any encoding; and tree-structured utility with explicit local factors.
  • Encoding-regime separation: Polynomial-time under the explicit-state encoding (polynomial in $|S|$). Under ETH, there exist succinctly encoded worst-case instances witnessed by a strengthened gadget construction (mechanized in Lean; see Appendix A) with $k^{*}=n$ for which SUFFICIENCY-CHECK requires $2^{\Omega(n)}$ time.
The tractable cases are stated with explicit encoding assumptions (Section 2.4). Together, these results answer the question “when is decision-relevant information identifiable efficiently?” within the stated regimes.
[…]
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 ($\sim$5,000 lines, 200+ theorems); complexity classifications follow by composition with standard results (see Appendix A).
Keywords: computational complexity, decision theory, polynomial hierarchy, tractability dichotomy, Lean 4
footnotetext: © 2026 Tristan Simas. This work is licensed under $C$ BY 4.0. License: https://creativecommons.org/licenses/by/4.0/
Zenodo DOI: 10.5281/zenodo.18140965
@@ 1 Introduction
Consider a decision problem with actions $A$ and states $S=X_{1}\times\cdots\times X_{n}$. A coordinate set $I\subseteq\{1,\ldots,n\}$ is sufficient if knowing only coordinates in $I$ determines the optimal action:
@@
This paper characterizes the efficient cases of coordinate sufficiency within the formal model:
Section 2.4 fixes the computational model and input encodings used for all complexity claims.
@@ 1.1 Landscape Summary
Each condition is stated with its encoding assumption. Outside these regimes, the general problem remains coNP-hard (Theorem 3.6).
@@
The lower-bound statement does not address intermediate regimes.
@@ 1.3 Machine-Checked Proofs
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 [6] ($\sim$5,000 lines, 200+ theorems). The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets. Complexity class membership (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follows by composition with standard complexity-theoretic results.
@@
We contribute (i) formal definitions of decision-theoretic sufficiency in Lean; (ii) machine-checked proofs of reduction correctness for the TAUTOLOGY and $\exists\forall$-SAT reductions; and (iii) a complete complexity landscape for coordinate sufficiency with explicit encoding assumptions. Prior work establishes hardness informally; we formalize the constructions.
@@ 1.4 Paper Structure
Section 2: foundations. Section 3: hardness proofs. Section 4: dichotomy. Section 5: tractable cases. Sections 7 and 8: corollaries and implications for practice. Section 9: related work. Appendix A: Lean listings.
@@ 2 Formal Foundations
We formalize decision problems with coordinate structure, sufficiency of coordinate sets, and the decision quotient, drawing on classical decision theory [12, 11].
@@ 2.2 Sufficiency and Relevance
Equivalently, the optimal action depends only on coordinates in $I$.
@@
Coordinate $i$ is relevant if it belongs to some minimal sufficient set.
@@ 2.3 The Decision Quotient
Coordinate set $I$ is sufficient if and only if $\text{DQ}_{I}(s)=|\text{Opt}(s)|/|A|$ for all $s\in S$.
@@ 2.4 Computational Model and Input Encoding
We fix the computational model used by the complexity claims.
@@
This succinct circuit encoding is the standard representation for decision problems in complexity theory; hardness is stated with respect to the input length of the circuit description [3]. An instance is encoded as:
@@
The input length is $L=|A|+\sum_{i}\log|X_{i}|+|C_{U}|$. Polynomial time and all complexity classes (coNP, $\Sigma_{2}^{P}$, ETH, W[2]) are measured in $L$. All hardness results in Section 3 use this encoding.

### v3

# Computational Complexity of Physical Information Sufficiency

@@ Abstract
We study physical information sufficiency as a decision-theoretic meta-problem. For $\mathcal{D}=(A,S,U)$ with factored state space $S=X_{1}\times\cdots\times X_{n}$, a coordinate set $I$ is sufficient when
$s_{I}=s^{\prime}_{I}\implies\text{Opt}(s)=\text{Opt}(s^{\prime}).$
This asks whether projected information preserves the full optimal-action correspondence.
Complexity landscape.
  • SUFFICIENCY-CHECK is coNP-complete; MINIMUM-SUFFICIENT-SET is coNP-complete; ANCHOR-SUFFICIENCY is $\Sigma_{2}^{\textsf{P}}$-complete.
  • Under explicit-state encoding, polynomial-time algorithms hold for bounded actions, separable utility, and tree-structured utility.
  • Under succinct encoding, hardness is regime-dependent: with ETH, there are worst-case families with $k^{*}=n$ requiring $2^{\Omega(n)}$ time.
  • Under query access, the finite-state core has worst-case Opt-oracle complexity $\Omega(|S|)$, with Boolean value-entry and state-batch refinements preserving the obstruction.
Physical grounding. The paper formalizes a physical-to-core encoding $E:\mathcal{P}\to\mathcal{D}$ and a transport rule: declared physical assumptions transfer to core assumptions, and core claims lift back to encoded physical instances. Encoded physical counterexamples induce core failures on the encoded slice. Discrete-time interface semantics (decision event = one tick) and budgeted thermodynamic lifts (bit lower bounds to energy/carbon lower bounds under declared constants) are formalized in the same assumption-typed framework. [C:P2.34,2.36,2.28,2.29,2.30,2.31,5.18,5.28;C2.35,5.22;T2.37;R:AR] (L: CT6,CT8,CT4,CT1,DQ548,DQ537,DQ539,DQ541,PBC1,CC65,DQ569.)
All theorem-level claims are machine-checked in Lean 4 (13391 lines, 597 theorem/lemma statements). Complexity-class completeness follows by composition with standard complexity results; regime-dependent and physical-transport consequences are proved as assumption-explicit closures.
Keywords: computational complexity, physical information, decision theory, polynomial hierarchy, formal verification
© 2026 Tristan Simas. This work is licensed under CC BY 4.0. License: https://creativecommons.org/licenses/by/4.0/
@@ 1 Introduction
Consider any decision problem $\mathcal{D}$ with actions $A$ and factored state space $S=X_{1}\times\cdots\times X_{n}$. A coordinate set $I\subseteq\{1,\ldots,n\}$ is sufficient if knowing only coordinates in $I$ determines the optimal action:
@@
This is coordinate sufficiency, the formalization of information sufficiency (whether available information determines the optimal action) over a factored state space.
In information-theoretic terms, this is a task-conditional coding problem: the object preserved by projection is the decision boundary $s\mapsto\text{Opt}(s)$, not the full source state. The same object appears in classical source-coding and fidelity viewpoints [57, 59, 14, 64], and in zero-error/confusability structure where exact recoverability depends on graph constraints rather than raw description length [58, 31, 38, 15]. The same separation also appears in model selection: short descriptions and certifiable adequacy are distinct [54, 24, 35, 8].
For physical instantiations, the same tuple has a direct substrate reading: coordinates are measurable state degrees of freedom, actions are physically realizable interventions or transitions, and the optimizer map is the decision boundary induced by the declared objective contract. The physical-to-core encoding discipline in Section 2.7 makes this bridge theorem-level rather than metaphorical.[C:P2.34,2.36,2.28,2.29,2.30,2.31;C2.35;T2.37;R:AR] (L: CT6,CT8,CT4,CT1,DQ548,DQ537,DQ539,DQ541.)
Compact claim-stamp notation.
We use compact provenance stamps of the form $C:\cdots;R:\cdots]$: claim references are grouped by type ($D/P/C/T/L/R$ for Definition/Proposition/Corollary/Theorem/Lemma/Remark), and regime tags are abbreviated (for example ARAR=active declared regime, DCDC=declared contract, CRCR=declared class/regime, QfQf=query-finite, QbQb=query-Boolean). Theorem-strength narrative statements are stamped and, where available, paired with compact Lean handle IDs; unstamped narrative remains explanatory or citation-only context.
SUFFICIENCY-CHECK, MINIMUM-SUFFICIENT-SET, and ANCHOR-SUFFICIENCY are meta-problems: each is parameterized by an arbitrary decision problem $\mathcal{D}$ and asks a structural question about $\mathcal{D}$’s coordinate space. The complexity results therefore apply to every problem with factored state space, not to a specific problem domain.[C:T4.6,4.7,4.9;R:S] (L: CC61,CC36,CC5.)
Remark 1.1 (Universality of the Decision-Problem Tuple).
The formal object of this paper is the decision-problem tuple $\mathcal{D}=(A,S,U)$: actions, factored states, utility. Within this model family, any problem requiring a choice among alternatives given structured information instantiates the tuple: the alternatives are $A$, the structured information is $S=X_{1}\times\cdots\times X_{n}$, and the objective is $U$. Thus the qualifier “decision problem” does not narrow scope inside the declared coordinate model. The meta-problem results below are universal over problems representable with this coordinate structure; the $(A,S,U)$ formalism supplies the language for stating the sufficiency predicate.
This paper characterizes the complexity landscape of these meta-problems within the formal model:
The paper keeps two layers explicit throughout: a substrate-neutral decision core (for complexity classification) and an assumption-typed physical instantiation layer (for transport to concrete systems). This separation prevents conflating representational hardness with physical assumptions while preserving a formal bridge between them.[C:P2.34;T2.37;R:AR] (L: CT6,CT4.)
Section 2.4 fixes the computational model and input encodings used for all complexity claims. Section 2.5 gives the model contract and regime tags used to type every strong claim. Completeness claims are class-internal: complete for the static sufficiency class fixed by C1–C4 and the declared access regimes. Adjacent objective classes that violate the bridge conditions are typed separately (Section 2.5).
@@
1.1 Hardness–Recovery Reading Map
Theorems in this paper are intended to be read as hardness/recovery pairs for the same fixed decision question:
  1. 1.
    Sufficiency checking: [S] coNP-complete in the general succinct regime (Theorem 4.6), paired with polynomial-time recovery under explicit-state and structured-access regimes (Theorem 6.1).
  2. 2.
    Minimal sufficient sets: [S] coNP-complete in general (Theorem 4.7), paired with the collapse criterion and relevance-computation recovery route (Theorems 4.14, 6.1).
  3. 3.
    Anchor sufficiency: $\Sigma_{2}^{\textsf{P}}$-complete in the general anchored formulation (Theorem 4.9); no general polynomial-time recovery condition is established in this model.
(L: CC60, CC36, CC34, CC4, CC71, S2P2.)
An operational criterion follows later from the same chain: once decision-site count exceeds the amortization threshold $n^{*}$, avoiding structural identification is more expensive than paying the one-time centralization cost (Theorem 9.6).[C:T9.6;R:LC] (L: HD3, HD21.)
@@ 1.2 Landscape Summary
Each condition is stated with its encoding assumption. Outside these regimes, the general problem remains coNP-hard (Theorem 4.6). [C:T4.6,6.1;R:E,S] (L: CC61,CC69,CC70,CC71.)
@@
  • In the intermediate query-access model: the finite-state core yields an Opt-oracle lower bound $\Omega(|S|)$ (Boolean instantiation: $\Omega(2^{n})$), and the mechanized Boolean interface refinements preserve the obstruction for value-entry and state-batch interfaces, with explicit subproblem-to-full transfer, weighted randomized robustness, and oracle-lattice transfer/strictness theorems (Propositions 5.35.16).
@@
The explicit ETH lower bound is still a succinct worst-case statement; Proposition 5.3 provides the finite-state query core and Propositions 5.75.16 provide theorem-level Boolean-interface refinements. [C:P5.3,5.7,5.8,5.11,5.15,5.16;R:Qf,Qb] (L: CC50,L9,L11,L8,L21,L20.)
@@ 1.3 Main Theorems
  6. 6.
    Theorem 2.37: physical-system bridge composition (law-objective semantics, behavior-preserving configuration reduction, and required-work lower bound) is proved as one bundled mechanized statement under declared encoding assumptions.
  7. 7.
    Propositions 2.282.31: the interface-time model is discrete, each decision event advances time by one unit, run-length equals elapsed interface time, and the unit-time law is substrate-invariant under the declared interface-consistency condition.
[C:T4.6,4.7,4.9,5.1,6.1,2.37;P2.28,2.29,2.30,2.31;R:E,S,S+ETH,AR] (L: CC61,CC36,CC5,CC16,CC69,CC70,CC71,CT4,DQ548,DQ537,DQ539,DQ541.)
1.4 Claim-Typing Theorems
  1. 1.
    Proposition 3.16: report admissibility is typed by certificate class ($\mathsf{ABSTAIN}$, $\mathsf{EXACT}$, $\mathsf{EPSILON}(\varepsilon)$).
  2. 2.
    Proposition 3.18: first-class evidence-object existence is equivalent to typed report admissibility.
  3. 3.
    Propositions 2.5 and 2.6: the $I=\emptyset$ query is exactly universal constancy of the decision boundary, and non-sufficiency is exactly existence of a counterexample pair.
  4. 4.
    Proposition 3.21: exact admissibility is equivalent to exact evidence existence.
  5. 5.
    Proposition 3.23 and Corollary 3.25: certified confidence is evidence-gated; for exact reports in no-competence regimes, certified confidence must be zero.
  6. 6.
    Proposition 3.24: under signal consistency, no evidence implies zero certified confidence.
  7. 7.
    Proposition 3.29 and Proposition 3.30: abstain may carry tentative guesses and self-reflected confidence, but self-confidence alone does not certify exactness.
  8. 8.
    Propositions 3.26, 3.27, 3.28: scalar step-count witnesses are always exact/falsifiable, while completion percentages are meaningful only under explicit declared bounds.
  9. 9.
    Theorem 3.56: outside declared carve-outs, admissible exact-claim policy requires explicit assumptions.
  10. 10.
    Proposition 3.61: under non-collapse, integrity and exact competence cannot both hold universally on hardness-blocked declared regimes.
  11. 11.
    Proposition 3.62 and Corollary 3.63: schema-level consequence for declared class/regime objects and exact-report inadmissibility.
[C:T3.56;P3.16,3.18,2.5,2.6,3.21,3.23,3.24,3.29,3.30,3.26,3.27,3.28,3.61,3.62;C3.25,3.63;R:DC,AR] (L: CC75,IC2,DP6,DP7,IC35,IC38,IC40,IC46,IC47,IC42,IC44,IC45,CC30,CC15,CC38.)
1.5 Regime Separation Theorems
  1. 1.
    Theorem 5.1: same-question separation between explicit-state and succinct regimes (with [S+ETH] hard-family lower bound).
  2. 2.
    Proposition 5.3: finite-state Opt-oracle lower bound in [Q_fin].
  3. 3.
    Propositions 5.75.16: Boolean-interface refinements and transfer/strictness in [Q_bool].
[C:T5.1;P5.3,5.7,5.11,5.8,5.15,5.16;R:E,S+ETH,Qf,Qb] (L: CC16,CC50,L9,L8,L11,L21,L20,L23.)
@@ 1.6 Machine-Checked Proofs
The reduction constructions and key equivalence theorems are machine-checked in Lean 4 [16, 17, 67, 21, 32] (13391 lines, 597 theorem/lemma statements). The formalization verifies that the TAUTOLOGY reduction correctly maps tautologies to sufficient coordinate sets. Complexity class membership (coNP-completeness, $\Sigma_{2}^{P}$-completeness) follows by composition with standard complexity-theoretic results.
@@
We contribute (i) formal definitions of decision-theoretic sufficiency in Lean; (ii) machine-checked proofs of reduction correctness for the TAUTOLOGY and $\exists\forall$-SAT reductions; (iii) a complete complexity landscape for coordinate sufficiency with explicit encoding assumptions; and (iv) a formal separation between structural complexity and representational hardness used to derive theorem-indexed engineering corollaries, including theorem-level typed coverage/completeness closures for the declared static class (Theorems 2.38, 3.4). Prior work establishes hardness informally; we formalize the constructions and their regime-typed consequences. [C:T2.38,3.4,4.6,4.9;R:CR,S] (L: CC14,CC76,CC61,CC5.)
@@ 1.7 Paper Structure
Section 2: decision-problem foundations and encoding model. Section 3: structural vs representational hardness; integrity vs competence. Section 4: hardness proofs. Section 5: regime separation. Section 6: tractable cases. Section 7: regime-conditional corollaries. Section 8: dominance theorems for hardness placement. Section 9: conservation law for decision-relevant coordinates. Section 10: related work. Appendix A: Lean listings.
Sections 79 are not applied commentary. Every claim in those sections is a formal corollary, regime-typed by the framework of Section 3 and indexed to a machine-checked theorem handle. They complete the complexity landscape by characterizing when exact minimization is rational and when it is not: the rationality matrix (Proposition 3.60), the dominance and conservation theorems (Theorems 8.16, 9.3), and the amortization threshold (Theorem 9.6) are all mechanized in Lean. [C:P3.60;T8.16,9.3,9.6;R:AR,LC] (L: IC27,HD1,HD4,HD3.)
@@ 2 Formal Foundations
We formalize decision problems with coordinate structure, sufficiency of coordinate sets, and the decision quotient, drawing on classical decision theory [55, 52]. The core definitions in this section are substrate-neutral: they specify the decision relation independently of implementation medium. Physical interpretation is introduced later through explicit encoding and assumption-transfer maps (Section 2.7).[C:D2.32;P2.34;T2.37;R:AR] (L: CT6,CT4.)
@@ 2.2 Sufficiency and Relevance
Equivalently, the optimal action depends only on coordinates in $I$. In this paper, this set-valued invariance of the full optimal-action correspondence is the primary decision-relevance target.
Proposition 2.5 (Empty-Set Sufficiency Equals Constant Decision Boundary).
For any decision problem,
$\text{sufficient}(\emptyset)\iff\forall s,s^{\prime}\in S,\ \text{Opt}(s)=\text{Opt}(s^{\prime}).$
Hence, the $I=\emptyset$ query is exactly a universal constancy check of the decision boundary. (L: DP6.)
Proposition 2.6 (Insufficiency Equals Counterexample Witness).
For any coordinate set $I$,
$\neg\text{sufficient}(I)\iff\exists s,s^{\prime}\in S:\ s_{I}=s^{\prime}_{I}\ \wedge\ \text{Opt}(s)\neq\text{Opt}(s^{\prime}).$
In particular,
$\neg\text{sufficient}(\emptyset)\iff\exists s,s^{\prime}\in S:\ \text{Opt}(s)\neq\text{Opt}(s^{\prime}).$
(L: DP7, DP8.)
[C:D2.4;P2.5,2.6;R:DM]
Definition 2.7 (Deterministic Selector).
A deterministic selector is a map $\sigma:2^{A}\to A$ that returns one action from an optimal-action set.
Definition 2.8 (Selector-Level Sufficiency).
For fixed selector $\sigma$, a coordinate set $I$ is selector-sufficient if
$\forall s,s^{\prime}\in S:\quad s_{I}=s^{\prime}_{I}\implies\sigma(\text{Opt}(s))=\sigma(\text{Opt}(s^{\prime})).$
Proposition 2.9 (Set-Level Target Dominates Selector Targets).
If $I$ is sufficient in the set-valued sense (Definition 2.4), then $I$ is selector-sufficient for every deterministic selector $\sigma$ (Definition 2.8). (L: DP5)
Proof.
If $I$ is sufficient, then $s_{I}=s^{\prime}_{I}$ implies $\text{Opt}(s)=\text{Opt}(s^{\prime})$. Applying any deterministic selector $\sigma$ to equal optimal-action sets yields equal selected actions.  $\blacksquare$
Definition 2.10 ($\varepsilon$-Optimal Set and $\varepsilon$-Sufficiency).
For $\varepsilon\geq 0$, define
$\text{Opt}_{\varepsilon}(s):=\{a\in A:\forall a^{\prime}\in A,\ U(a^{\prime},s)\leq U(a,s)+\varepsilon\}.$
Coordinate set $I$ is $\varepsilon$-sufficient if
$\forall s,s^{\prime}\in S:\ s_{I}=s^{\prime}_{I}\implies\text{Opt}_{\varepsilon}(s)=\text{Opt}_{\varepsilon}(s^{\prime}).$
Proposition 2.11 (Zero-$\varepsilon$ Reduction).
Set-valued sufficiency is exactly the $\varepsilon=0$ case:
[…]
(L: DP1, DP4)
Proposition 2.12 (Selector-Level Separation Witness).
The converse of Proposition 2.9 fails in general: there exists a decision problem, selector, and coordinate set $I$ such that $I$ is selector-sufficient but not sufficient in the full set-valued sense. (L: CC54.)
@@
Coordinate $i$ is relevant if there exist states that differ only at coordinate $i$ and induce different optimal action sets:
$i\text{ is relevant}\iff\exists s,s^{\prime}\in S:\;\Big(\forall j\neq i,\;s_{j}=s^{\prime}_{j}\Big)\ \wedge\ \text{Opt}(s)\neq\text{Opt}(s^{\prime}).$
Proposition 2.15 (Minimal-Set/Relevance Equivalence).
For any minimal sufficient set $I$ and any coordinate $i$:
$i\in I\iff i\text{ is relevant}.$
Hence every minimal sufficient set is exactly the relevant-coordinate set. (L: DP2, DP3.)
Proof.
The “only if” direction follows by minimality: if $i\in I$ were irrelevant, removing $i$ would preserve sufficiency, contradicting minimality. The “if” direction follows from sufficiency: every sufficient set must contain each relevant coordinate.  $\blacksquare$
Definition 2.16 (Exact Relevance Identifiability).
For a decision problem $\mathcal{D}$ and candidate coordinate set $I$, we say $I$ is exactly relevance-identifying if
$\forall i\in\{1,\ldots,n\}:\quad i\in I\iff i\text{ is relevant for }\mathcal{D}.$
Equivalently, $I$ is exactly relevance-identifying iff $I$ equals the full relevant-coordinate set.
@@ 2.3 The Decision Quotient
[C:P2.20;R:H=succinct-hard] Coordinate set $I$ is sufficient if and only if $\text{DQ}_{I}(s)=|\text{Opt}(s)|/|A|$ for all $s\in S$. (L: CC62, CC63)
@@ 2.4 Computational Model and Input Encoding
We fix the computational model used by the complexity claims. In physical deployments, these encodings correspond to different access conditions on the same underlying system: explicit tables correspond to fully exposed state/utility structure, succinct encodings correspond to compressed generative descriptions, and query regimes correspond to constrained measurement interfaces.[C:T5.1;P5.3,2.34;R:E,S+ETH,Qf,AR] (L: CC16,CC50,CT6.)
@@
This succinct circuit encoding is the standard representation for decision problems in complexity theory; hardness is stated with respect to the input length of the circuit description [5, 47]. An instance is encoded as:
@@
The input length is $L=|A|+\sum_{i}\log|X_{i}|+|C_{U}|$. Polynomial time and all complexity classes (coNP, $\Sigma_{2}^{P}$, ETH) are measured in $L$. All hardness results in Section 4 use this encoding.
@@
Query-access regime (intermediate black-box access).
The solver is given oracle access to decision information at queried states (e.g., $\text{Opt}(s)$, or $U(a,s)$ with $\text{Opt}(s)$ reconstructed from finitely many value queries). Complexity is measured by oracle-query count (optionally paired with per-query evaluation cost). This separates representational access from full-table availability and from succinct-circuit input length.